OpenStudy (anonymous):

Find the exact value of the expression: csc^-1 ((-2√3)/3) Explain please!

OpenStudy (anonymous):

i think a good first step would be to recognize \(-\frac{2\sqrt{3}}{3}\) as \(-\frac{2}{\sqrt{3}}\)

OpenStudy (anonymous):

and also recalling that \(\csc(x)\) is the reciprocal of \(\sin(x)\)

OpenStudy (anonymous):

that way, you can ask "do i know an angle whose sine is \(-\frac{\sqrt{3}}{2}\) ? if you do, you are done

OpenStudy (anonymous):



OpenStudy (anonymous):

Oy I didn't think of it that way, so it's π/3?

OpenStudy (anonymous):

this is not tangent

OpenStudy (anonymous):

actually it is \(-\frac{\pi}{3}\) because it is negative

OpenStudy (anonymous):

Oh right! (:

OpenStudy (anonymous):

picture works nice for problems like \[\tan(\cos^{-1}(x))\] i.e.compositions

OpenStudy (anonymous):

for these you just sort of have to know the answer

OpenStudy (anonymous):

Thanks again c: